OpenStudy (anonymous):

Erica Tomlinson's credit card company determines her minimum monthly payment by adding all new interest to 1% of the outstanding principal. The credit card company charges an interest rate of 0.03456% per day. On October 15, Erica uses her credit card to purchase a custom set of wheels and tires for her car for $2200. She makes no other purchases or new interest during October. On November 1, instead of making the minimum payment, Erica makes a payment of $500. Assuming there are no additional charges or cash advances, determine Erica's minimum payment due on December 1.

OpenStudy (anonymous):

@dumbcow hay buddy help this guy i have no idea about this one :)

OpenStudy (dumbcow):

my thinking is since it was purchased in October, no interest will be charged until November... she paid 500 so outstanding balance on Nov. 1 is 1700 interest accrues for 30 days new balance = 1700(1.0003456)^30 = 1717.71 interest = 17.71 min payment = 17.71 + 17 = 34.71

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

ty dude :)

OpenStudy (dumbcow):

no problem...btw advice always pay more than min payment :)